# Spurious reconstruction from brain activity

### v3

@@ Results
Finally, we conducted a rigorous evaluation of the generator component, which typically consists of diffusion models in text-guided reconstruction methods. To ensure that a visual image reconstruction method has the potential to faithfully reproduce an individual’s perceived visual experiences, it is crucial that the method can recover the original images with a high degree of perceptual similarity when the neural translation from brain activity to latent features is perfect. However, it has been unclear whether recent text-guided reconstruction methods meet this fundamental requirement. To address this question, we performed a recovery check analysis by reconstructing images using the true latent features of target images. Instead of using latent features translated from brain activity, we directly input the latent features derived from the target images into the generator.
@@
Text-guided methods produced images semantically similar to targets but not perceptually similar (Fig 7A), while the iCNN method yielded results that closely resembled the actual target images. These findings suggest that text-guided reconstruction methods may prioritize semantic similarity over perceptual accuracy. In contrast, the iCNN method appears to have a superior ability to capture and replicate original visual content, indicating potential advantages in preserving fine-grained visual details.
@@
Fig 7: Recovery check. (A) Reconstruction from the true latent features. The leftmost column displays the original images. Subsequent columns show reconstruction results from the StableDiffusionReconstruction, Brain Diffuser, and iCNN methods. (B) Component-wise reconstruction of the Brain-Diffuser method. The leftmost column shows the original images. The following columns present reconstruction results using individual components of the Brain-Diffuser method (VDVAE, CLIP text, and CLIP vision features). Each row represents a different test image. The reconstruction results reveal the upper bound of reconstruction performance for each method and highlight potential limitations in the latent feature representations or generative processes.
@@
To investigate the recovery performance further, we conducted a recovery check on each latent feature of the Brain-Diffuser method (Fig 7B). Interestingly, reconstructions from VDVAE features, which are used for generating initial images in the Brain-Diffuser, exhibited a high degree of similarity to the target images. However, the images generated by CLIP features through the diffusion models showed significant deviations from the original targets. These findings suggest that text-guided reconstruction methods may not be well-suited for visual image reconstruction tasks, as they fail to faithfully recover the original visual images. Instead, they tend to create images based on their latent features, such as CLIP features, which can lead to a phenomenon known as “hallucination” in the field of generative AIs. Hallucination refers to an output that appears plausible but is actually incorrect or misleading, raising concerns about the reliability and accuracy of the model [59]. Text-guided reconstruction methods seem to prioritize generating semantically similar images rather than faithfully reconstructing the visual content perceived by the individual.
@@
The above findings may provide an explanation for why the text-guided reconstruction methods performed well only on the NSD dataset (Fig 2A). The results of the case study demonstrated that the text-guided reconstruction methods struggle to reconstruct (Fig 2B) or identify (Fig 6) test samples that lie beyond the distribution of the training set. Such limitations suggest that these methods lack true generalization capabilities and are unable to accurately reconstruct novel visual experiences that differ significantly from the examples they were trained on. Moreover, even when the test samples belonged to the same distribution as the training set, the translators of the text-guided reconstruction methods had difficulty correctly identifying those test samples (Fig 5). This observation indicates that the translators may not have learned a sufficiently robust and generalizable mapping between brain activity patterns and the corresponding latent features, further limiting their ability to faithfully reconstruct the perceived visual experiences.
@@
This case study also revealed that the NSD test stimuli are highly similar to the training set, with a significant overlap in their visual and semantic features (Fig 4). Given this similarity, the impressive reconstruction results achieved by the recent text-guided reconstruction methods on the NSD dataset should not be interpreted as evidence of zero-shot reconstruction capabilities. Instead, a more plausible interpretation is that these methods primarily function as a combination of “classification” and “hallucination.”
@@
In this context, the translator component of the text-guided reconstruction methods functions primarily as a classifier, predicting categorical semantic information present in the training phase rather than capturing the fine-grained details of the visual experience. This limitation may explain why convincing reconstructions can be obtained even from random brain data through post-hoc selection (Fig 3B). Due to the limited variety in the outputs generated by the reconstruction models, repeated trials and post-hoc selection can eventually find images that are semantically and visually similar to the target stimulus. The apparent plausibility and semantic consistency of these generated images can be attributed to the capabilities of the diffusion model, which learns to generate realistic-looking images based on semantic information. While these images may seem convincing at first glance, they do not accurately reflect the specific visual experience of the individual. This phenomenon of hallucination raises serious concerns about the reliability and validity of the text-guided reconstruction methods when evaluated on the NSD dataset. Although we evaluated only two reconstruction methods in this case study, it is important to recognize that any reconstruction methods evaluated only by NSD [46, 48, 60] can potentially have similar classification and hallucination problems.
Formal analysis and simulation
@@
Building on the issues identified in our case study, we examine these challenges in a more general context. We have identified several issues with these text-guided reconstruction methods and the dataset, including the cluster structure of CLIP latent features, the lack of diversity in NSD, and the misspecification of the latent representation for image reconstruction. These issues resulted in the inability of diffusion models to faithfully recover the original images from their latent features. However, it is crucial to recognize that the findings of the case study are not merely specific to CLIP, NSD, or diffusion models. Instead, these issues likely reflect more fundamental problems that can arise in the development and evaluation of brain decoding and visual image reconstruction methods. Thus in this section, we extend the problems identified in the case study into formal analyses and simulations in generalized settings, aiming to provide a more comprehensive understanding of the factors that contribute to the limitations of current reconstruction methods and explore strategies for mitigating these issues.
Output dimension collapse
@@
Multivariate linear regression models are widely used in constructing decoding and encoding models of the brain. These models are often regarded as capable of independent and compositional predictions for each target, as they create separate regression models for individual targets without sharing the weights between them. However, this expectation is not generally true. This is particularly evident when input variables are shared [33, 35, 34, 44, 45]. In this section, we demonstrate this assertion by examining the problem of predicting multiple targets using linear regression models with shared inputs.
@@
Next, we consider a scenario where the diversity of the target features is small. This situation can arise when the feature space exhibits a clustered structure and the training data lacks sufficient diversity, as observed in the case study with the CLIP text features and the NSD dataset (Fig 4A). When the training features have limited diversity, the predicted values from brain activity, which are represented as linear combinations of these target features, also become constrained. Consequently, the prediction from brain data to target features effectively becomes a projection onto a low-dimensional subspace formed by the training data.
@@
Fig 8: Demonstration of output dimension collapse in feature prediction. The left panel shows the distribution of source brain activity data in the first two dimensions of the original high-dimensional space. The right panel displays the distribution of target latent features projected onto the first two principal component (PC) dimensions. A linear ridge regression model is trained using the training data (gray points in the left and right panels), in which the output features in the training data are constrained to two clusters. When presented with test data from the brain space (left panel), the model generates predictions (orange points) in the latent feature space (right panel). This visualization demonstrates how the predicted features are constrained to the subspace defined by the training data, highlighting the limitation in generalizing beyond the training data distribution.
@@
To illustrate this phenomenon, we simulated teacher-student learning, a framework where a “teacher” model generates data based on certain underlying rules, and a “student” model is trained to learn or approximate those rules by observing the generated data. We examined the distribution of predicted values from the student linear regression model trained on clustered features generated by the teacher model. We generated clustered features by sampling from a Gaussian mixture distribution in a high-dimensional space. The corresponding brain activity samples were generated from the latent feature samples multiplied by teacher weight and adding observation noise (See Materials and Methods “Simulation with clustered data”). Student weights were obtained by training a linear regression model to predict the clustered feature values from the corresponding brain data. We then projected the randomly generated brain samples into latent feature space by learned regression weight. To visualize the high-dimensional predicted patterns effectively, we projected the predicted features onto the PCA spaces derived from the training features and showed the first two dimensions.
@@
The simulation results clearly demonstrate the impact of clustered features on the predicted values (Fig 8). The trained linear regression model projects arbitrary brain data onto the subspace defined by the latent features in the training set, resulting in predicted values that are confined to the vicinity of the training clusters. This observation highlights the limitation of training linear regression models with clustered features as prediction targets: the trained model’s predictions are inherently constrained by the diversity and structure of the training data.
@@
We refer to this phenomenon as “output dimension collapse,” where the model’s predictions become confined to a limited subspace of the output feature space. It has important implications for the generalization capability of linear regression models in the context of brain decoding and visual image reconstruction. When training data lack diversity and form distinct clusters in the feature space, the translator overly adapts to the subspace formed by the training data, regardless of the potential of the latent feature space. Consequently, the translator’s outputs become confined to patterns similar to those in the training set, severely limiting the model’s ability to predict novel or out-of-distribution samples.
@@
Output dimension collapse may explain why plausible reconstructions were obtained even from random brain data by merely generating images several times in the case study (Fig 3). The lack of semantic diversity in NSD causes the translator to adapt only to the feature patterns of the training set, restricting its outputs to the subspace formed by the training data. As a result, convincing images could be found even from random data through questionable post-hoc selection.
It is crucial to understand that output dimension collapse may not inherently be a limitation of linear regression models. This phenomenon can manifest across various regression models, including nonlinear approaches such as multi-layer perceptrons (MLPs) [46] and transformers [48]. Nonlinear models may be more susceptible to output dimension collapse due to their enhanced flexibility in adapting to training data.
@@
It is also important to recognize that the mathematical formulation in Eqs 1–3 assumes that all input variables are shared across target variables. If each target is predicted from a distinct set of input variables through feature (voxel) selection, the predictions can become more independent, potentially mitigating output dimension collapse. This approach has been utilized in the field since its early days, with techniques such as sparse voxel selection and modular modeling [28, 29, 61, 37].
@@
Simulation with clustered features: What makes prediction compositional?
@@
The case study revealed that the NSD exhibits limited diversity (Fig 4) and poses difficulties for zero-shot prediction (Fig 5 and Fig 6D). These observations suggest that the translator of CLIP features suffers from output dimension collapse due to the lack of semantic diversity in the NSD. To explore potential strategies for mitigating output dimension collapse and achieving flexible predictions, we conducted simulation analyses using clustered features to assess generalization performance beyond the training set.
@@
As in the previous section, our simulation involved teacher-student learning. We first generated feature data ${\mathbf{y}}\in\mathbb{R}^{D}$ then made the input brain data ${\mathbf{x}}\in\mathbb{R}^{D}$ by translating ${\mathbf{y}}$ with the teacher weight and added observation noise. To simulate a situation where the dataset has cluster structures and to control diversity effectively, the training feature vector $\mathbf{y}\in\mathbb{R}^{D}$ was generated from a $D$-dimensional Gaussian mixture (Fig 9A).
@@
We trained a ridge regression model on large training data samples and obtained the student weight. To simulate the situation where the trained model encounters clusters that are not available at the training phase, we used two types of test samples, in-distribution and out-of-distribution test samples: in-distribution test samples were generated from one of the clusters used in the training set, whereas out-of-distribution (OOD) samples were generated from the novel cluster that is not included in the training set. For these two types of predicted features, we calculated the cluster identification accuracy (Fig 6C) by using $C+1$ cluster centers: $C$ centers from the training set and one cluster center of the OOD test set.
@@
Fig 9: Simulation analysis of predicting features with a cluster structure (A) Illustration of target latent features setting. The latent features were generated from Gaussian mixture distributions. $\sigma_{\mathrm{intra}}$ is a parameter that controls the scale of the samples within a cluster. $\sigma_{\mathrm{inter}}$ controls the scale between the cluster centers. In-distribution test samples are generated from one of the clusters in the training set. Out-of-distribution (OOD) test samples are generated from a novel cluster not included in the training set. (B) Cluster identification accuracy. The $x$-axis represents the number of clusters used in the training, and the $y$-axis shows the cluster identification performance. The green line represents the results of in-distribution test samples, and the orange line shows the results for OOD samples. The dashed curve indicates the chance level. (C) Scenarios for achieving generalizability with sufficient data diversity. The upper illustration shows the training data covering the whole latent feature space, requiring an exponential order relative to the feature dimension. The lower one shows that the training data covers only the effective axes of the latent feature space, leading to a linear order relative to the feature dimension. (D) Relation between latent feature dimension and the sufficient number of clusters for generalization. The $x$-axis represents the dimension of the target latent features. The $y$-axis shows the number of clusters that achieve above $0.5$ cluster identification accuracy. The three lines indicate the results from different ratios of $\sigma_{\mathrm{intra}}^{2}/\sigma_{\mathrm{inter}}^{2}$. (E) Example of successful prediction beyond the training distribution in a 2D output feature space. Dark green points represent the in-distribution target features, with their corresponding light green points showing the model’s predictions. Dark orange points denote the OOD target features, with light orange points indicating their predictions. Despite the OOD target features (dark orange) not being included in the training clusters, they are accurately predicted (light orange), demonstrating the model’s generalization ability. This successful generalization highlights the power of compositional prediction, where the model effectively combines individual dimensions to predict novel feature combinations outside the training distribution.
@@
We first examined zero-shot prediction performance for different numbers of training clusters (Fig 9B). We fixed the feature dimension $D$ and the cluster variance ratio constant. While the cluster identification performances of in-distribution test samples were perfect, the performances of OOD test samples showed different patterns depending on the number of clusters in the training data. When the number of training clusters was small, the cluster identification accuracy was 00%. This result indicates that the behavior of the translator became more similar to a classifier, making it difficult to generalize beyond the training set, as observed in NSD cases (see also Fig 6D). On the other hand, as the number of clusters in the training data increased, it became possible to identify the novel clusters, achieving the same performance as in-distribution test samples. This observation indicates the importance of the diversity of the training dataset. We also emphasize that large numbers of training samples do not necessarily address the problem. All of these results were obtained with a sufficiently large amount of training data, and the number of training clusters was varied while keeping the amount of training data fixed. Also, we observed qualitatively similar results in increasing the data diversity by controlling the cluster variance ratio and keeping the number of dimension and training clusters constant (S4 Fig.A).
@@
Next, we investigate how diverse the training data needs to be to ensure sufficient generalization. There are two possible scenarios for diversifying training samples: either by densely sampling the entire target feature space so that there are no remaining gaps (Fig 9C; top) or by uniformly sampling to the extent that it covers the entire dimension of the target feature space (Fig 9C; bottom). The former scenario requires an exponentially larger number of samples/clusters relative to the dimension, whereas the latter only requires up to a linear order. We sought to reveal which of these two scenarios was more likely to be true by varying the dimensions of the feature space and identifying the number of clusters required for generalization (S4 Fig. B for identification accuracy in each condition). Here, we defined the number of clusters required for generalization as the point at which the identification accuracy of OOD samples exceeds $50$%. The relationship between the dimension of the feature space and the number of clusters necessary for generalization appears to be linear (Fig 9D). This finding suggests that achieving generalization does not necessarily require an exponentially large diversity that fills the entire feature space. Instead, it suffices to have a number of clusters that cover the adequate dimensions within the target feature space. Although obtaining a large amount of brain data is hard work, it is important for a dataset to contain sufficiently diverse stimuli covering the effective dimensions of the target feature space to achieve zero-shot prediction.
@@
We also confirmed this phenomenon with a simple and transparent example ($D=2$, Fig 9E). The training data covers sufficient axes in the target feature space, enabling the prediction of locations not present in the training set. Based on this low-dimensional intuition, we argue that successful zero-shot prediction requires that training data leads to representations that can serve as a basis for spanning the target feature space. Leveraging such bases effectively enables the model to predict novel samples by predicting each basis and combining them. This compositional representation, spanning the target feature space, is crucial for zero-shot prediction [62] and reconstructing arbitrary visual images from limited brain data.
Caveat with evaluation by pairwise identification

### v5

@@ 2.1.4 Failed recovery of a stimulus from its latent features
Finally, we conducted a rigorous evaluation of the generator component, which typically consists of diffusion models in text-guided reconstruction methods. To ensure that a visual image reconstruction method has the potential to faithfully reproduce an individual’s perceived visual experiences, it is crucial that the method can recover the original images with a high degree of perceptual similarity when the neural translation from brain activity to latent features is perfect. However, it has remained unclear whether recent text-guided reconstruction methods meet this fundamental requirement. To address this question, we performed a recovery check analysis by reconstructing images using the true latent features of target images. Instead of using latent features translated from brain activity, we directly input the latent features derived from the target images into the generator.
@@
Text-guided methods produced images semantically similar to targets but not perceptually similar (Fig. 7A), while the iCNN method yielded results that closely resembled the actual target images. These findings suggest that text-guided reconstruction methods may prioritize semantic similarity over perceptual accuracy. In contrast, the iCNN method appears to have a superior ability to capture and replicate original visual content, indicating potential advantages in preserving fine-grained visual details.
@@
Fig. 7: Recovery check. (A) Reconstruction from the true latent features. The leftmost column displays the original images. Subsequent columns show reconstruction results from the StableDiffusionReconstruction, Brain Diffuser, and iCNN methods. (B) Component-wise reconstruction of the Brain-Diffuser method. The leftmost column shows the original images. The following columns present reconstruction results using individual components of the Brain-Diffuser method (VDVAE, CLIP text, and CLIP vision features). Each row represents a different test image. The reconstruction results indicate the upper bound of reconstruction performance for each method and highlight potential limitations in the latent feature representations or generative processes.
@@
To investigate the recovery performance further, we conducted a recovery check on each latent feature of the Brain-Diffuser method (Fig. 7B). Interestingly, reconstructions from VDVAE features, which are used for generating initial images in the Brain-Diffuser, exhibited a high degree of similarity to the target images. However, the images generated by CLIP features through the diffusion models showed significant deviations from the original targets. These findings suggest that text-guided reconstruction methods may not be well-suited for visual image reconstruction tasks, as they fail to faithfully recover the original visual images. Instead, they tend to create images based on their semantic features, such as CLIP features, which can lead to a phenomenon known as “hallucination” in the field of generative AIs. Hallucination refers to an output that appears plausible but is actually incorrect or misleading, raising concerns about the reliability and accuracy of the model (Rawte et al., 2023). Text-guided reconstruction methods seem to prioritize generating semantically similar images rather than faithfully reconstructing the visual content perceived by the individual (see Fig. A8 for the recovery with MindEye2’s generators).
@@
The above findings may provide an explanation for why the text-guided reconstruction methods performed well only on the NSD dataset (Fig. 2A). The results of the case study demonstrated that the text-guided reconstruction methods struggle to reconstruct (Fig. 2B) or identify (Fig. 6D) test samples that lie beyond the distribution of the training set. Such limitations suggest that these methods lack true generalization capabilities and are unable to accurately reconstruct novel visual experiences that differ significantly from the examples they were trained on. Moreover, even when the test samples belonged to the same distribution as the training set, the translators had difficulty correctly identifying those test samples (Fig. 5). This observation indicates that the translators may not have learned a sufficiently robust and generalizable mapping between brain activity patterns and the corresponding latent features, further limiting their ability to faithfully reconstruct the perceived visual experiences.
@@
This case study also revealed that the NSD test stimuli are highly similar to the training set, with a significant overlap in their visual and semantic features (Fig. 4). Given this similarity, the impressive reconstruction results achieved by the recent text-guided reconstruction methods on the NSD dataset should not be interpreted as evidence of zero-shot reconstruction capabilities. Instead, a more plausible interpretation is that these methods primarily function as a combination of “classification” and “hallucination.”
@@
In this context, the CLIP feature translator in the text-guided reconstruction methods functions primarily as a classifier, predicting categorical semantic information present in the training phase rather than capturing the fine-grained details of the visual experience. This limitation may explain why convincing reconstructions can be obtained even from random brain data through post-hoc selection (Fig. 3BC). Due to the limited variety in the outputs generated by the reconstruction models, repeated trials and post-hoc selection can eventually find images that are semantically and visually similar to the target stimulus. The apparent plausibility and semantic consistency of these generated images can be attributed to the capabilities of the diffusion model, which learns to generate realistic-looking images based on semantic information. While these images may seem convincing at first glance, they do not accurately reflect the specific visual experience of the individual. This phenomenon of hallucination raises serious concerns about the reliability and validity of the text-guided reconstruction methods when evaluated on the NSD dataset. Although we particularly evaluated three reconstruction methods in this case study, it is important to recognize that any reconstruction methods evaluated only by NSD (Scotti et al., 2023; Quan et al., 2024) can potentially have similar classification and hallucination problems due to the limited diversity of the NSD. Furthermore, recent highly realistic reconstruction methods leveraging generative AI models (Chen et al., 2023; Bai et al., 2024; Benchetrit et al., 2024) should also be carefully validated to ensure their performance is not overly dependent on dataset biases (Fig. A5).
2.2 Formal analysis and simulation
@@
Building on the issues identified in our case study, we examine these challenges in a more general context. We have identified several issues with these text-guided reconstruction methods and the dataset, including the cluster structure of CLIP latent features, the lack of diversity in the NSD, and the misspecification of the latent representation for image reconstruction. These issues resulted in the inability of diffusion models to faithfully recover the original images from their latent features. However, it is crucial to recognize that the findings of the case study are not merely specific to CLIP, NSD, or diffusion models. Instead, these issues likely reflect more fundamental problems that can arise in the development and evaluation of brain decoding and visual image reconstruction methods. Thus, in this section, we extend the problems identified in the case study into formal analyses and simulations in generalized settings, aiming to provide a more comprehensive understanding of the factors that contribute to the limitations of current reconstruction methods and explore strategies for mitigating these issues.
2.2.1 Output dimension collapse
@@
Multivariate linear regression models are widely used in constructing decoding and encoding models of the brain. These models are often regarded as capable of independent and compositional predictions for each target, as they create separate regression models for individual targets without sharing the weights between them. However, this expectation is not generally true. This is particularly evident when input variables are shared (Seeliger et al., 2018; Ozcelik et al., 2022; Mozafari et al., 2020; Takagi and Nishimoto, 2023a; Ozcelik and VanRullen, 2023). In this section, we demonstrate this assertion by examining the problem of predicting multiple targets using linear regression models with shared inputs.
@@
Next, we consider a scenario where the diversity of the target features is small. This situation can arise when the feature space exhibits a clustered structure and the training data lacks sufficient diversity, as observed in the case study with the CLIP text features and the NSD dataset (Fig. 4A). When the training features have limited diversity, the predicted values from brain activity, which are represented as linear combinations of these target features, also become constrained. Consequently, the prediction from brain data to target features effectively becomes a projection onto a low-dimensional subspace formed by the training data.
@@
Fig. 8: Demonstration of output dimension collapse in feature prediction. The left panel shows the distribution of source brain activity data in the first two dimensions of the original high-dimensional space. The right panel displays the distribution of target latent features projected onto the first two principal component (PC) dimensions. A linear ridge regression model is trained using the training data (gray points in the left and right panels), in which the output features in the training data are constrained to two clusters. When presented with test data from the brain space (left panel), the model generates predictions (orange points) in the latent feature space (right panel). This visualization demonstrates how the predicted features are constrained to the subspace defined by the training data, highlighting the limitation in generalizing beyond the training data distribution.
@@
To illustrate this phenomenon, we simulated teacher-student learning, a framework where a “teacher” model generates data based on certain underlying rules, and a “student” model is trained to learn or approximate those rules by observing the generated data. We examined the distribution of predicted values from the student linear regression model trained on clustered features generated by the teacher model. We generated clustered features by sampling from a Gaussian mixture distribution in a high-dimensional space. The corresponding brain activity samples were generated from the latent feature samples multiplied by teacher weight and adding observation noise (see Methods “Simulation with clustered data”). Student weights were obtained by training a linear regression model to predict the clustered feature values from the corresponding brain data. We then projected the randomly generated brain samples into the latent feature space using the learned regression weights. To visualize the high-dimensional predicted patterns effectively, we projected the predicted features onto the PCA spaces derived from the training features and showed the first two dimensions.
@@
The simulation results clearly demonstrate the impact of clustered features on the predicted values (Fig. 8). The trained linear regression model projects arbitrary brain data onto the subspace defined by the latent features in the training set, resulting in predicted values that are confined to the vicinity of the training clusters. This observation highlights the limitation of training linear regression models with clustered features as prediction targets: the trained model’s predictions are inherently constrained by the diversity and structure of the training data.
@@
We refer to this phenomenon as “output dimension (or domain) collapse,” where the model’s predictions become confined to a limited subspace (or subdomain) of the output feature space. It has important implications for the generalization capability of linear regression models in the context of brain decoding and visual image reconstruction. When training data lack diversity and form distinct clusters in the feature space, the translator overly adapts to the subspace formed by the training data, regardless of the potential of the latent feature space. Consequently, the translator’s outputs become confined to patterns similar to those in the training set, severely limiting the model’s ability to predict novel or out-of-distribution samples.
@@
Output dimension collapse may explain why plausible reconstructions were obtained even from random brain data by merely generating images several times in the case study (Fig. 3). The lack of semantic diversity in the NSD causes the translator to adapt only to the feature patterns of the training set, restricting its outputs to the subspace formed by the training data. As a result, convincing images could be found even from random data through questionable post-hoc selection.
It should be noted that this phenomenon is not inherently limited to linear regression models; it can occur in various multivariate regression models, including multi-layer neural networks. In fact, when we replaced the nonlinear translator in the MindEye2 reconstruction with a linear translator, the reconstructions’ bias toward object categories in the training set was substantially reduced (Fig. A9). This observation suggests that nonlinear models are more susceptible to the collapse of the output domain due to their greater flexibility in fitting to training data.
@@
It is also important to recognize that the mathematical formulation in Eqs. 1–3 assumes that all input variables are shared across target variables. If each target is predicted from a distinct set of input variables through feature (voxel) selection, the predictions can become more independent, potentially mitigating output dimension collapse. This approach has been utilized in the field since its early days, with techniques such as sparse voxel selection and modular modeling (Miyawaki et al., 2008; Yamashita et al., 2008; Fujiwara et al., 2013; Shen et al., 2019b).
@@
2.2.2 Simulation with clustered features: What makes prediction compositional?
@@
The case study revealed that the NSD exhibits limited diversity (Fig. 4) and poses difficulties for zero-shot prediction (Fig. 5 and Fig. 6D). These observations suggest that the translator of CLIP features suffers from output dimension collapse due to the lack of semantic diversity in the NSD. To explore potential strategies for mitigating output dimension collapse and achieving flexible predictions, we conducted simulation analyses using clustered features to assess generalization performance beyond the training set.
@@
As in the previous section, our simulation involved teacher-student learning. We first generated feature data ${\mathbf{y}}\in\mathbb{R}^{D}$ then made the input brain data ${\mathbf{x}}\in\mathbb{R}^{D}$ by translating ${\mathbf{y}}$ with the teacher weight and added observation noise. To simulate a situation where the dataset has cluster structures and to control diversity effectively, the training feature vector $\mathbf{y}\in\mathbb{R}^{D}$ was generated from a $D$-dimensional Gaussian mixture (Fig. 9A).
@@
We trained a ridge regression model on large training data samples and obtained the student weight. To simulate the situation where the trained model encounters clusters that are not available at the training phase, we used two types of test samples, in-distribution and out-of-distribution test samples: in-distribution test samples were generated from one of the clusters used in the training set, whereas out-of-distribution (OOD) samples were generated from the novel cluster that is not included in the training set. For these two types of predicted features, we calculated the cluster identification accuracy (Fig. 6C) by using $C+1$ cluster centers: $C$ centers from the training set and one cluster center of the OOD test set.
@@
Fig. 9: Simulation analysis of predicting cluster-structured features. (A) Illustration of target latent features. The latent features were generated from Gaussian mixture distributions, with $\sigma_{\mathrm{intra}}$ controlling within-cluster spread and $\sigma_{\mathrm{inter}}$ controlling inter-cluster scaling. In-distribution test samples are generated from training clusters while Out-of-distribution (OOD) test samples come from novel clusters. (B) Cluster identification accuracy for various numbers of training clusters. The $x$-axis represents the number of training clusters, and the $y$-axis shows the cluster identification performance. The green and orange lines indicate results for in-distribution and OOD samples, respectively. The dashed curve indicates the chance level. (C) Scenarios for achieving generalizability with sufficient data diversity. The upper illustration shows the training data covering the whole latent feature space, requiring an exponential order relative to the feature dimension. The lower one shows that the training data covers only the effective axes of the latent feature space, leading to a linear order relative to the feature dimension. (D) Sufficient number of clusters for generalization as a function of latent feature dimension. The $x$-axis represents the dimension of the target latent features. The $y$-axis shows the number of clusters achieving above $0.5$ cluster identification accuracy, with curves for different $\sigma_{\mathrm{intra}}^{2}/\sigma_{\mathrm{inter}}^{2}$ ratios. (E) Example of successful prediction beyond the training distribution in a 2D output feature space. In-distribution and OOD target features are depicted in dark green and dark orange, and the model’s predictions of these features are depicted in light green and light orange, respectively. Despite the OOD target features (dark orange) not being included in the training clusters, they are accurately predicted (light orange), demonstrating the model’s generalization ability by combining learned feature dimensions.
@@
We first examined zero-shot prediction performance for different numbers of training clusters (Fig. 9B). We fixed the feature dimension $D$ and the cluster variance ratio constant. While the cluster identification performances of in-distribution test samples were perfect, the performances of OOD test samples showed different patterns depending on the number of clusters in the training data. When the number of training clusters was small, the cluster identification accuracy was 00%. This result indicates that the behavior of the translator became more similar to that of a classifier, making it difficult to generalize beyond the training set, as observed in the NSD cases (see Fig. 6D). On the other hand, as the number of clusters in the training data increased, it became possible to identify the novel clusters, achieving the same performance as in-distribution test samples. This observation indicates the importance of the diversity of the training dataset. We also emphasize that large numbers of training samples do not necessarily address the problem. All of these results were obtained with a sufficiently large amount of training data, and the number of training clusters was varied while keeping the amount of training data fixed. Also, we observed qualitatively similar results in increasing the data diversity by controlling the cluster variance ratio while keeping the number of dimensions and training clusters constant (Fig. A10A).
@@
Next, we investigate how diverse the training data needs to be to ensure sufficient generalization. There are two possible scenarios for diversifying training samples: either by densely sampling the entire target feature space so that there are no remaining gaps (Fig. 9C; top) or by uniformly sampling to the extent that it covers the entire dimension of the target feature space (Fig. 9C; bottom). The former scenario requires an exponentially larger number of samples/clusters relative to the dimension, whereas the latter only requires up to a linear order. We sought to reveal which of these two scenarios was more likely to be true by varying the dimensions of the feature space and identifying the number of clusters required for generalization (Fig. A10B for identification accuracy in each condition). Here, we defined the number of clusters required for generalization as the point at which the identification accuracy of OOD samples exceeds $50$%. The relationship between the dimension of the feature space and the number of clusters necessary for generalization appears to be linear (Fig. 9D). This finding suggests that achieving generalization does not necessarily require an exponentially large diversity that fills the entire feature space. Instead, it suffices to have a number of clusters that cover the adequate dimensions within the target feature space. Although obtaining a large amount of brain data is hard work, it is important for a dataset to contain sufficiently diverse stimuli covering the effective dimensions of the target feature space to achieve zero-shot prediction.
@@
We also confirmed this phenomenon with a simple and transparent example ($D=2$, Fig. 9E). The training data covers sufficient axes in the target feature space, enabling the prediction of locations not present in the training set. Based on this low-dimensional intuition, we argue that successful zero-shot prediction requires training data to leads representations that can serve as a basis for spanning the target feature space. Leveraging such bases effectively enables the model to predict novel samples by predicting each basis and combining them. This compositional representation, spanning the target feature space, is crucial for zero-shot prediction (Schug et al., 2024) and reconstructing arbitrary visual images from limited brain data.
While our simulations varied the number and spread of clusters to examine the role of training diversity, we assumed that cluster centers were uniformly distributed across the latent feature space. However, this assumption may not accurately reflect the characteristics of actual datasets such as NSD, where cluster centers themselves can potentially be biased in the entire visual space (i.e., focusing on natural scenes in MSCOCO). Even with a large number of clusters, generalization to unseen images remains challenging if they occupy only a limited region of the visual space. These considerations emphasize that true diversity requires careful control over both the number and the spatial distribution of training clusters or images.
2.2.3 Caveat with evaluation by pairwise identification